OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. M(0, 6), N(6, 0)

OpenStudy (anonymous):

eq of line is \[\left( y-y1 \right)=\frac{ y2-y1 }{x2-x1}\left( x-x1 \right)\] simplify and then write in the form ax+by+c=0

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

yw